DOC

Specification

By Rick Adams,2014-01-20 03:51
7 views 0
Specification

    ?????

     COMMISSIONING REPORT

    STND1. SATURDAY / SUNDAY 21 22 APR (0800-2100 / 0800-1600)

    Returned to the vessel in Seattle in order to complete the outstanding work

    rdprior to the departure of the vessel on 23 April. The work done was as follows:

    -

    a) Joystick modification (Point 36) - The modification to remove the signal isolators was completed. This was done to the supplied drawing. Some of the

    wire numbers were modified. Information on the changes will be forwarded

    to Mike Bartlett on return to UK. The modification did not initially work

    because the signal polarity was the opposite to that required. It could not be

    simply swapped round due to the removal of the isolators. The problem was

    resolved by modifying the GEM80 software to reverse the incoming polarity of

    the joystick signal.

    b) Crompton meter scales (Point 28) - New scales were fitted on the meters

    as follows: -

    LHS 0 20 x1000 lb

    RHS 0 40 x 1000 lb

    c) OC1 Drive (Point 26) - Fitted replacement 63A fuse. N.B. Also missing the

    GSL1000 trip indicator. A spare trip indicator will be taken from OC2 drive since the OC1 drive is to be used in the tests next week.

    d) Totco equipment (Points 22, 23 and 24) The Aft and RHS Totco units were

    not powered up on arrival. This proved to be as a result of USCG work on the

    Totco system. The fault was as a result of the reversal of the 12V power supply

    cabling at the main Totco units. This was rectified and now all the units are

    powered up. The cabling problem came about because a lot of the wiring in

    the Totco boxes is un-marked. The alarms page for the system also indicated

    that the four off sheave serial links had failed. This was found to be as a result

    of loss of mains power to the server (although there was no alarm for this). This

    was corrected and left winch 4 sheave serial link failure only. N.B. This is

    because the winch 4 unit has been removed and sent for repair. To

    summarise: - Sheaves 1 3 are now fully functional. The tests next week require winch 4 sheave in operation (Drum 2 is to be used).

    This means that the winch 3 Totco equipment needs to be moved to meet

    the winch 4 cabling (Action USCG).

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 1 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    N.B. When the T/C drive was operated from the local (downstairs) joystick

    (with winch 4 available / sheave 4 link fault) the drive ran for several seconds

    then stopped and the winch 3 available lamp came on instead of winch 4

    but the drive didn’t run. Moving the joystick back to zero re-selected winch 4

    and the operation could be repeated. This needs some investigation.

    e) Trawl / Core winch labelling (Point 31) The label supplied is the same as

    the original. The label should read: -

    0 Trawl Core 9/16"

    1 Trawl Core 3/4"

    The dimensions between hole centres should be as below in order to avoid

    having to re-drill the door: -

    Horizontal 3.7/8"

    Vertical 4.25"

    The winch 3 / winch 4 label was fitted (door re-drilled).

    f) Trawl / Winch drive (Point 27) - Re-fitted 4A fuse in FS6. g) Trawl / Winch drive (Point 18) - Fitted the new thermostat. h) Other points The aluminium boxes appear to have been removed from

    the ship. It is unknown where they have been taken too. No news on the

    Pittsburgh spares that were due to arrive on Friday. Unable to locate the

    winch spares which have also been moved.

    The e-mail address for me should be as follows: -

    JQPUBLICHEALY@HEALY.USCG.MIL Please try e-mailing the Totco analogue output set-up information again.

    G Pyatt to confirm a delivery address for San Francisco tomorrow (Jerry).

    i) Some points were raised by the ‘Scientists’ (N.B. these are to be discussed further since Glen considers that these facilities are either already available or

    not necessary. They should perhaps be looked into): -

    - Can the new logging system be used to look at data (e.g. tension) in real-

    time?

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 2 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    - They are suggesting that it may be a good idea to have a tension meter

    (or another Totco unit) on the bridge. This would also involve a new cable

    to the bridge.

    - They are suggesting adding a recorder (real time) onto the Totco

    analogue output to record tension as a real time trend. The main concern

    would be to ensure the recorder has a high enough input impedance.

    RD2. MONDAY 23 APR (0800-2000)

    Completing the outstanding points as follows: -

    a) I checked the operation of the tension meters with a battery. The meter

    pointers showed no signs of stiction.

    b) In conjunction with the USCG the No.3 Totco unit equipment was moved

    into the No.4 position. The unit is presently reading a relatively high tension figure at zero tension and obviously needs re-calibrating. The RS232-485

    converter on No.3 Totco is still faulty. It was also found that the analogue

    output module had suffered two bent pins on installation. Attempts were

    made to straighten these pins but one broke off. A replacement module is

    now required for the unit.

    c) I investigated the control changeover problem on winch 3. This is as a

    result of the modification put in to select the drum number. The drum number

    is being set to zero in the drive as soon as the drive runs. This causes the

    control to changeover. The problem only seems to occur when control is at

    the local (downstairs) control panel. At the aft control panel the drive runs as normal. The drive software was modified to correct this error and the drive

    now works correctly locally.

    d) The correct T/C wire selection label drilling details are as follows: -

    Horizontal 3.7/8"

    Vertical 4.1/4"

    e) Advised Mike Bartlett regarding the various queries and the e-mail address

    for communications.

    TH3. TUESDAY 24 APRIL (0800 2000)

    Preparatory operational tests were conducted on winches 1 and 4 prior to

    there use tomorrow. The key items of note from these tests were as follows: -

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 3 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    a) Calibration of the Totco units Winch 1 Totco unit was found to be

    displaying a much faster speed than the winch was running. Closer investigation showed that a Trawl / Core winch Totco unit had been installed in the winch 1 position. The decision was made to re-calibrate the payout speed and tension on both of the winch 1 and 4 units (N.B. winch 2 appears to be calibrated correctly). The speed calibration for winch 1 was copied from winch 2 (winch 4 was already calibrated correctly). A two position calibration test was done in order to calibrate tension. This was completed successfully although the reading accuracy was found to be reduced at lower tensions, sometimes resulting in negative tension readings.

    b) During the Totco unit calibration tests the analogue outputs were set up. These did seem to work as per the manual. Finalised settings for al the Totco units were as follows: -

0lb 47.88

    20000lb 83.96

    c) The Trawl / Core winch 4 stopped twice during a cast operation. The drive had lost Control On and this condition could not be reset. Investigations showed that Control On request could be written on at the drive. Loss of Control On can only occur by loss of either: -

    A32.6 LCN closed contact

    G300.9 Software Control On request

    L9932.0 Bridge A available (DC Link pre-charged)

    In order to progress the problem I did the following: -

    1. Inhibited the FiP failure trip (Set G4 = 0) 2. Put in a software trap to catch the appropriate tables on loss of G10.6

    (Retained LCN).

    This to be further investigated tomorrow while dredging.

    TH4. WEDNESDAY 25 APR (0800-2000 / 2200-0300)

    The first dredge operation showed up two problems on the trawl /core winch drive as follows: -

    a) DC Link Overvoltage trips After some investigations regarding the level of

    the link voltage (131%) and the setting of the Overvoltage trip (150%) the

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 4 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    parameters were left at the original settings. To alleviate the trips the decision was made to revert back to the original speed control stability figures. The following parameters were changed: -

    L2121 = 4.5s (was 6.5s)

    L2122 = 2.5s (was 1s)

    Although there was the suggestion that the above figures could be switched based on the scope (e.g. scope >50m: de-gain) the sheave may not be

    on-line. This would mean that the scope figure seen at the drive would be zero and consequently the drive would be on the high gain figures even though it could have an actual scope >50m.

    The drive appeared to control satisfactorily with the dredge both in and out of the water with the lower gain figures so the parameters were left at these settings ahead of further discussion on the implementation of the gain change if still thought necessary.

    b) The drive was still shutting down occasionally (Control On Off). This

    effectively eliminated a FiP fault as a possibility (G4 set to 0) and the software trap indicated that A32.6 was causing the problem. No problems could be seen on either the wiring of A32.6 or the wiring of LCN and LCN pilot coils. The input was wire linked out for test purposes and no further trips occurred. The status of both a) and b) is to be further reviewed tomorrow.

    The actual trawl was ultimately successful with basalt rocks being retrieved from the seabed approximately 3.5km below the ship.

    During the evening an earth fault was reported on OC1 winch drive. This proved to be as a result of several resistors in a harmonic filter circuit exploding. The resistors involved are R4-6 (3 sets of 3 x parallel resistors). There was evidence of serious overheating on all these particular resistors and their associated wiring. Further examination of the other two drives showed that their resistors were in a similar condition to those in OC1. There were three spare resistors. These were used to improve the Trawl/Core winch circuit to an acceptable condition. The other two winch drive filter circuits cannot be used so it is necessary to check if the drive can be ran without the filter circuit. The current through the resistors and the voltage across them was measured and found to be 8A and 15V respectively. All the information was passed back for a review on component values and ratings. Replacement resistors will be required for all three drives when the values and ratings are confirmed.

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 5 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    TH5. THURSDAY 26 APR (0800-2000)

    The Trawl/Core winch was used without further problems.

    M Bartlett confirmed that the winch drives could be ran without the filter circuits for a short period and so OC1 drive was made available and ran satisfactorily. N.B. two main isolator fuses were found blown in OC1 drive. Spares were taken from the OC2 drive which now needs two new fuses (Bussmann 170M1569).

    Some voltage harmonics measurements were carried out on the main supply to OC1 and the trawl/core winch. These gave fairly low values of THD (<6% all conditions) and confirmed that the resistors appear to be simply under-rated. Two things of note from the operations on the winches as follows: - a) The GEM80 appears to lose Totco data for short periods occasionally. This manifests itself as the scope light flashing and the logged data going to zero. This only happens on the trawl/core winch.

    b) The joystick speed inputs are not symmetrical and do not have a 0-100% range. I have queried the accuracy of the resistors.

    TH6. FRIDAY 27 APR (0800-2400)

    I had an overnight report of the trawl/core winch running away (paying out) twice during reel-in. This also happened during the first dredge today. This appears to be because of the ‘drive stopping’ problem. This is being manifested in this different way now that A32.6 is linked out. The fact that it is still happening proves that the A32.6 circuit is working OK. I have removed this wire link and added a temporary LCN latching contact across RLY12 (pre-charged). This should eliminate the pre-charge circuit. I also had an overvoltage trip while paying out and have increased the overvoltage trip point to 155%.

    Put in modifications that act to reduce the speed ramp under certain conditions. These provide for the following: -

    Fine speed control 6s

    Full speed control 20s

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 6 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    TH7. SATURDAY 28 APR (0800-2300)

    As the OC1 winch drive was energised during preparations for a cast the

    main FiP A and B hubs tripped out. Investigations into this followed on the

    OC1 drive as follows: -

    FiP B connection only into the drive Alarm at ECC (FiP system B failure)

    FiP A connection only into the dive Appeared to be OK for five minutes then

    shut down the FiP A system.

    The showed the FiP board was faulty and so it was replaced with a spare with

    no further problems.

    This is a very serious problem since steerage was lost from the bridge (ship

    under emergency control from ECC).

    Work continued on the trawl/core winch run-away/stopping problems. It was

    demonstrated that RLY12 (DC Link charged) seemed to be sometimes

    operating when the drive was running and hence causing LCN to de-

    energise. If the RLY12 contact is latched in, in the LCN circuit, then run-away

    occurs. If not then the drive stops with LCN open in the fault condition.

    Did several other peripheral tests as follows: -

    Checked the encoder speed feedback on an oscilloscope OK

    Replaced the PIB with a board from OC2 No effect. N.B. USCG have no

    spare board.

    Modified the trawl/core drive code to change the control for the drum

    selection of the emergency brakes.

    TH8. SUNDAY 29 APR (0800 2200) During preparations to run the trawl/core winch it was found that the winch

    control GEM80 had halted. Closer examination showed that the GEM80 had

    halted with a 703 fault (program CRC mismatch code corruption) and the

    GEMLAN D board had lost its healthy lamp. The GEM80 was recovered by re-

    loading the program, but he GEMLAN D card would not recover. This was

    replaced with a spare board. Following this the system recovered.

    The speed control of the trawl/core drive was stabilised with wire deployed.

    A bandwidth of approximately 2 radians was set.

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 7 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    It was noted that the speed feedback responded quite differently between

    the step and the joystick. This was referred back to Rugby.

    Tests showed that the drive appears to have a speed related problem above

    approximately 65mpm.

    TH9. MONDAY 30 APR (0800 0200)

    Since it was not now possible to easily do dredges due to water depth, the

    tests were moved onto drum 1 (winch 3) with banded wire.

    Drum 1 was ran up and was found to stop in the same manner as happened

    with Drum 2 during the winch dredges (at approx. 65 - 70 mpm).

    Various tests were done as follows: -

    A step response was done on the DC Link The response was OK

    The speed ramp rate was reduced to 10%/sec and the drive still found to trip

    if the speed stability was pushed up to the original figures or the Overvoltage

    trip point was reduced to the original figure or if the ramp rate was increased

    (in fact the ramp rate had to be reduced temporarily to 8%/sec as a trip

    occurred at 10%/sec).

    The DC Link set-point was reduced to 125% from 131% (U/V trip = 75% O/V trip

    155%) and running tests continued unable to cause the drive to stop.

    Completed a dredge, running at speeds of up to top speed in both directions

    as well as at around the normal 65-75mpm speed and also trying speed

    changes unable to cause the drive to stop.

    It was noticed during the tests that the current limit (L2124) appeared to

    reduce to approx. 15% at around 70% speed with the reduced DC Link figure.

    It was also shown that this effect happened with 131% DC Link voltage in the

    form of a dip. If the load on the drive was increased slightly at his point then

    the current limit returned to around 80-90%. The reason for the reduced

    current limit is L1109.0 switching the current clamp (G70.14) The L1109.0

    table is Vq at maximum limit. This effect has been referred back to Rugby.

    During the course of the tests during the day (no drives running but OC1 and

    Trawl/Core winch with Control On) there was a further FiP disturbance that

    caused both FiP hubs to drop out and lose steering from the bridge. This

    disturbance seemed to simply recover and required no action from me.

    The three brake lifted micro-switches were examined (Point 19) and two were

    adjusted to work. One other is broken. The mechanical fixing of the switches

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 8 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

    ?????

    are poor and an improved bracket is required to fix them properly. The

    system was left with the original one micro-switch only in circuit as previously.

    I have asked the Interoceans representative to forward a piece of fibre-optic

    cable to USCG (Glen) to allow the end of rope detector to be tested as a

    complete system.

    There was an incident where the joystick was left off the zero point following a

    drive stop with loss of control available. The drive was reset and immediately

    started. The code should have protection for this to check.

    ST10. TUESDAY 1 MAY (0800 - )

    Return to port at San Fransisco approx. 1100.

     USCG Cutter Healy A Customer : Issue : SWC Systems Upgrade 01 May ‘01 Project : Date : Commissioning Diary 3 G Pyatt Document : Author : CB04261 9 Ref : Page : ?2001 ALSTOM Controls Ltd ALSTOM Power Conversion Ltd

Report this document

For any questions or suggestions please email
cust-service@docsford.com